OpenStudy (samigupta8):

The vector (i*a.b)i +(j*a.b)j+(k*a.b)k is equal to Here i,j,k are unit vectors in x ,y and z directions...and a.b represent the scalar product of two vectors a and b...

OpenStudy (samigupta8):

* symbol stands for cross product

OpenStudy (samigupta8):

@priyar

OpenStudy (samigupta8):

Options are a*b b*a a b

OpenStudy (priyar):

what is the cross product of a scalar and a vector?

OpenStudy (samigupta8):

Exactly !..i also thought the same

OpenStudy (priyar):

ya..it can't be that.. * must denote simple multiplication...not cross prod.

OpenStudy (samigupta8):

R u sure?? I mean how can we do then

OpenStudy (priyar):

coz a.b = scalar ..for ex: say 2 so i*a.b = 2i

OpenStudy (samigupta8):

N its multiplication with i again will give us??

OpenStudy (samigupta8):

It will give us scalar ryt?

OpenStudy (samigupta8):

I mean if we assume your example only then 2i multiplied with i gives 2

OpenStudy (priyar):

ya but that is dot product.. is (2i)i = 2i.i ?

OpenStudy (priyar):

seeing the options i can say that they have intended * to be cross product.. so forget what we did abv.

OpenStudy (samigupta8):

Actually it cannot possibly be that simple product as we proposed it earlier

OpenStudy (priyar):

i think the question is wrong coz cross product of a vector and a scalar doesn't make any sense..like i x 2 =?? doesn't have any physical meaning..

OpenStudy (samigupta8):

Well, I don't think so bcoz it's not even given as bonus in ans key

OpenStudy (priyar):

bonus?

OpenStudy (samigupta8):

If any question is wromg or it"z options doesn't match with the correct ans then ans key gives bonus remark on it

OpenStudy (michele_laino):

we can use the theory of determinants, so we can write: \[\Large i \times a = \det \left( {\begin{array}{*{20}{c}} i&j&k \\ 1&0&0 \\ {{a_x}}&{{a_y}}&{{a_z}} \end{array}} \right) = \left( {\begin{array}{*{20}{c}} {0,}&{ - {a_z},}&{{a_y}} \end{array}} \right)\] therefore, we get: \[\Large \left( {i \times a} \right) \cdot b = {a_y}{b_z} - {a_z}{b_y} = {\left( {a \times b} \right)_x}\] Similarly for other components Then the the resultant vector is: \[\Large {\left( {a \times b} \right)_x}i + {\left( {a \times b} \right)_y}j + {\left( {a \times b} \right)_z}k = a \times b\]

OpenStudy (samigupta8):

Gr8 sir .....wow!!!

OpenStudy (michele_laino):

thanks!! :D

OpenStudy (priyar):

yeah! why didn't we think of that.. just wrong grouping of terms!